# A Reinforcement Learning Approach for RIS-aided Fair Communications

### v1

@@ I Introduction
In this paper, we address these challenges by proposing a novel framework that integrates DRL with fairness-aware mechanisms to enhance RIS-aided communication systems. We present experimental findings and simulation outcomes, and we make our code openly available to encourage further research in this area. Based on the existing literature, our approach uses DRL algorithms such as Deep Deterministic Policy Gradient (DDPG) and Twin Delayed Deep Deterministic Policy Gradient (TD3). This choice allows for a more effective comparison with the studies upon which our work is founded, enabling us to concentrate on the additional parameters that configure our decision-making agent. Thus, we refrain from utilizing advanced algorithms such as Soft-Actor Critic (SAC) [8] and Proximal Policy Optimization (PPO) [9], to ensure a consistent progression between our work and the articles we are based on. Our contribution includes a proposal of new reward functions to ensure equitable service distribution among users while maintaining high system performance and a discussion about the experimental results.
@@ IV Fairness Improvements
The objective is to develop a system that maximizes the legitimate data rate transmitted to UE, ensuring that each user receives a minimum usable data rate. Additionally, we aim to equitably distribute any surplus data rate among users. The goal is to prevent any single user or group from monopolizing the majority of the bandwidth, ensuring that others are not left with minimal resources. This optimization process is non-linear, making it difficult to solve using traditional methods. The presence of multiple constraints further complicates the underlying problem. Next, we summarize the reward functions of the two main approaches surveyed in Section II.
@@ IV-A Common Terms
The reward function presented in [12] forms the basis of our formulation, emphasizing pure capacity across all users. For each user $i$, the secrecy rate $R_{i}$ is defined as:
@@
where $[x]^{+}$ corresponds to $\max(0,x)$. The Sum of Squared Rewards (SSR), final reward, is given by:
@@
The reward function presented in [13] takes into account the Quality of Service (QoS) requirements associated to the users. The reward proposed in [13] is only for the downlink scenario and with the presence of eavesdroppers. We therefore adapt it to keep the terms that are relevant for our duplex case without any eavesdroppers:
$R_{QoS}$ $=\sum_{i=1}^{k}R_{i}-\sum_{i=1}^{k}\mu_{1}(p_{i}^{D}+p_{i}^{U})$ (5)
with $\mu_{1}$ a constant used to balance the utility and
@@ IV-D Extended Reward Function
We propose two extensions to the reward function proposed in Peng et al.  [12] by adding, as well a QoS metric. These rewards are intended to reflect, for a designated user $i$, wether or not the system is currently ensuring the minimum capacity constraint and the resultant amount of data rate received.
IV-D1 Fairness and Quality of Service
@@
$R_{\text{fairness}}=\frac{\left(\sum_{i=1}^{k}R_{\text{Q}}^{i}\right)^{2}}{k\ \cdot\sum_{i=1}^{k}{R_{\text{Q}}^{i}}^{2}}$ (9)
which corresponds to the Jain’s fairness index [14] for the $R_{\text{Q}}^{i}$ rewards for all users $i$.
Notice that $R_{\text{fairness}}$ takes into account the fairness of the obtained rewards. It helps to better (and fairly) distribute the capacity of the system so one user is not obtaining an overly large proportion of the data rate, leading to deteriorated service for the other users.
@@
where $\alpha\in[0,1]$ is a hyperparameter that sets the tradeoff between pure QoS and pure fairness reward. We further refer to this reward function by using the Fairness & Quality of Service (FQoS) acronym for the experiments curves.
IV-D2 Fairness and Quality of Service Score
@@ V-A Implemented Algorithms
In addition to the DDPG algorithm [15] used in [12], we also implement and use TD3 [4] to better handle the instability of the environment and mitigate the residuals errors and overestimation for which DDPG is prone to. These two single-agent actor critic architectures are summarized in Figure 2. The state given as an input to the DRL system is kept unchanged from Peng et al.  [12]. The algorithms are implemented without any modifications from their respective original paper [15] [16] and only the reward functions have been modified to match our desired goals.
We evaluate and illustrate the impact of the reward systems proposed in Section IV, w.r.t. system capabilities and fairness. We present the results of conducting multiple training sequences, each utilizing different options for both algorithms and reward functions.
The environment remains consistent across all runs: the BS, positioned at coordinates $(0,0)$, communicates with the RIS located at $(20,100)$ in both directions. 2 stationary UE units are randomly distributed at the beginning of each episode within a square area defined by the $x$-axis range $[125,200]$ and the $y$-axis range $[25,100]$
@@
We outline only the most critical parameters to facilitate comprehension and highlight significant changes between sequences. The learning rates are $\delta_{a}=5\cdot 10^{-3}$ for the actor and $\delta_{c}=1\cdot 10^{-3}$ for the critic. The target actor and critics employ soft updates with $\tau=1\cdot 10^{-4}$. The actor performs a gradient descent update every two steps, while the critic updates every step. We assume a replay buffer containing $200001$ samples with a batch size of $2048$. Our experiments primarily involve multiple parallel instances of the environment for roll-outs, gathering diverse experiences to populate the replay buffer. This approach allows the optimization process to sample from a broader set of experiences, minimizing correlation.
We employ $15$ parallel environments, each initialized with different internal seeds, to conduct the roll-outs. All experiments use the same master random seed to ensure reproducibility. For each experiment, we present the evolution of the average reward stored in the replay buffer during training. Since the replay buffer operates on a first-in, first-out basis, this average serves as a reliable indicator of the system’s learning progress from a reward perspective. The objective of the DRL system is to maximize the obtained reward, regardless of the chosen reward function.
To compare the different systems derived from various reward functions, we also plot the evolution of rewards obtained with each possible reward function. The reward function used for the optimization process is termed the decisive reward function, while the others are referred to as informative. By examining the three curves for each system, we can effectively compare their performances and behavior using a consistent standard. Additional metrics recorded and stored during the training process, along with the precise code used to generate these metrics.
@@ V-B Downlink and Uplink Power Patterns
Figure 3 illustrates the downlink and uplink power patterns generated using our simulation framework. This figure presents an example of beam patterns for both downlink and uplink scenarios in a duplex communication system with a single user equipment. The yellow curves represent the beam power patterns transmitted by the BS toward the RIS for the user in the downlink scenario. The purple curves depict the beamforming power patterns at the RIS level using the projected polar profile. It is directed toward either the user or the BS, depending on whether it is downlink or uplink communication. The green rectangle represents an obstacle blocking the initial signal from the BS, highlighting the necessity of using the RIS.
@@ V-C Simulation Results
Figure 4 illustrates the average reward in the replay buffer at time $t$. This metric helps us assess whether the system is learning and improving in terms of its decisive reward functions. Given that the replay buffer operates on a first-in-first-out basis, it provides a convenient means to analyze the learning progress of each system. However, this graph does not allow us to determine which reward system performs best, as the three possible reward functions operate on different scales. Nonetheless, it enables us to track the evolution of a specific system.
To facilitate a better comparison of algorithm performance, Figure 5 depicts the evolution of reward in Equation  (4) during training for all systems, even though it is only informative for some. We observe that while systems using this reward as their decisive reward excel at maximizing it, other systems—particularly those using the FQoS reward function defined in Equation (10) can achieve results approximately $90\%$ as effective as those attained by systems using reward in Equation (4). To easily compare the fairness of each system, we provide Figure 6 that is displaying the evolution of the Jain fairness index through the whole training process. This index is computed using Equation (9) for the decisive reward chosen per user.
@@ V-D Discussion
A notable trend observable in our results is that systems utilizing reward function denoted in Equation (4) quickly achieves high performance from a reward perspective (around a nominal value of $6$ bps/Hz, in terms of spectral efficiency) and then experience a slight decline, yet remain within a high reward range for this particular function. Systems employing other reward functions exhibit similar behavior, suggesting that our systems are learning effectively, although most progress appears rapid due to the frequency of our measurements.
Figure 6 shows that systems employing the reward function defined in Equation (4) achieve a mean fairness of $0.6$ for both DDPG and TD3 algorithms. When the reward function defined in Equation (10) is used (as the decisive reward function), then the mean fairness rises to $0.8$. Given that the Jain fairness index here ranges from $\frac{1}{n}$ (worst case, where one user dominates) to $1$ (best case, where resources are perfectly split), with an $n$ value of $2$, the reward function defined in Equation (10) enhances fairness by nearly $40\%$, compared to the original reward function in Equation (4).
Figure 6: Mean Jain fairness index for each episode. Smoothed version with a rolling window of size $10$.
@@ VI Conclusion
We explored the concept of fair communications with respect to DRL-enabled RIS environments. We assumed that RISs must ensure that UE units receive their signals with adequate strength, without other UE being deprived of service due to insufficient power. We addressed such a problem and explored the fairness properties of previous work. We proposed a novel method that aims at obtaining an efficient and fair duplex DRL-RIS system, for multiple legitimate UE units. We reported experimental work and simulation results. We have also released our code and datasets to foster further research in the topic. Perspectives for future work include a more exhaustive analysis of complex environments by adding mobility patterns to the users and increasing the scale of the whole system, i.e number of users or RIS elements.

### v2

@@ I Introduction
In this paper, we address these challenges by proposing a novel framework that integrates DRL with fairness-aware mechanisms to enhance RIS-aided communication systems. We present experimental findings and simulation outcomes, and we make our code openly available to encourage further research in this area. Based on the existing literature, our approach uses DRL algorithms such as Deep Deterministic Policy Gradient (DDPG) and Twin Delayed Deep Deterministic Policy Gradient (TD3). This choice allows for a more effective comparison with the studies upon which our work is founded, enabling us to concentrate on the additional parameters that configure our decision-making agent. Thus, we refrain from utilizing state of the art algorithms such as Soft-Actor Critic (SAC) [8] and Proximal Policy Optimization (PPO) [9], to ensure a consistent progression between our work and the articles we are based on. Our contribution includes a proposal of new reward functions to ensure equitable service distribution among users while maintaining high system performance and a discussion about the experimental results.
@@
$R_{i}$ Secrecy Rate for user $i$
$\alpha$ Balance parameter between fairness and capacity
$C$ Sum Secrecy Rate
@@ IV Fairness Improvements
The objective is to develop a system that maximizes the legitimate data rate transmitted to UE, ensuring that each user receives a minimum usable data rate. As stated above we want to provide a fair service among all users despite the non-linearity of our optimization problem. Next, we summarize the reward functions of the two main approaches surveyed in Section II.
@@ IV-A Common Terms
The reward function presented in [12] forms the basis of our formulation, emphasizing pure capacity across all users. For each user $i$, the Secrecy Rate (SR) $R_{i}$ is defined as:
@@
where $[x]^{+}$ corresponds to $\max(0,x)$. The Sum Secrecy Rate (SSR), final reward, is given by:
@@
The reward function presented in [13] takes into account the Quality of Service (QoS) requirements associated to the users. The reward proposed in [13] is only for the downlink scenario and with the presence of eavesdroppers. We therefore adapt it to keep the terms that are relevant for our duplex case without any eavesdroppers. We propose the following reward that we denote as QoS reward:
$R_{QoS}$ $=\sum_{i=1}^{k}R_{i}-\sum_{i=1}^{k}\mu(p_{i}^{D}+p_{i}^{U})$ (5)
with $\mu$ a constant used to balance the combined performances and the individual constraints for each user.
@@ IV-D Extended Reward Function
We propose an extension to the reward function proposed in Peng et al.  [12] by adding a QoS and a fairness incentived term. This reward is intended to reflect, for a designated user $i$, wether or not the system is currently ensuring the minimum capacity constraint and the resultant amount of data rate received.
@@
$R_{\text{fairness}}=\frac{\left(\sum_{i=1}^{k}R_{i}\right)^{2}}{k\ \cdot\sum_{i=1}^{k}{R_{i}}^{2}}$ (9)
which corresponds to the Jain Fairness Index (JFI) [14] for the SR $R_{i}$ for all users $i$.
Notice that $R_{\text{fairness}}$ takes into account the fairness regarding SR and SSR rewards because it is the effective data rate obtained in reality. It is supposed to help to better (and fairly) distribute the capacity of the system so one user is not obtaining an overly large proportion of the data rate, leading to deteriorated service for the other users.
@@
where $\alpha\in[0,1]$ is a hyperparameter that sets the tradeoff between pure QoS and pure fairness reward. We further refer to this reward function by using the Fairness & Quality of Service (FQoS) acronym.
@@ V-A Implemented Algorithms
In addition to the DDPG algorithm [15] used in [12], we also implement and use TD3 [4] to better handle the instability of the environment and mitigate the residuals errors and overestimation for which DDPG is prone to. These two single-agent actor critic architectures are summarized in Figure 2. The state given as an input to the DRL system is kept unchanged from Peng et al.  [12]. The algorithms are implemented without any modifications from their respective original paper [15, 16] and only the reward functions have been modified to match our desired goals.
We evaluate and illustrate the impact of the reward systems proposed in Section IV in terms of system capabilities and fairness. We present the results of conducting multiple training sequences, each utilizing different options for both algorithms and reward functions.
The environment remains consistent across all runs: the BS, positioned at coordinates $(0,0)$, communicates with the RIS located at $(20,100)$ in both directions. 2 stationary UE units are randomly distributed at the beginning of each episode within a square area defined by the $x$-axis range $[125,200]$ and the $y$-axis range $[25,100]$ and with a precision of 0.1 for each coordinate. Superposition is made impossible.
@@
We outline only the most critical parameters to ease comprehension and highlight significant changes between sequences. The learning rates are $\delta_{a}=5\cdot 10^{-3}$ for the actor and $\delta_{c}=1\cdot 10^{-3}$ for the critic. The target actor and critics employ soft updates with $\tau=5\cdot 10^{-4}$. The actor performs a gradient descent update every two steps, while the critic updates every step. We assume a replay buffer containing $200001$ samples with a batch size of $2048$. Our experiments primarily involve multiple parallel instances of the environment for roll-outs, gathering diverse experiences to populate the replay buffer. This approach allows the optimization process to sample from a broader set of experiences, minimizing correlation.
We utilize 15 parallel environments, each initialized with distinct internal seeds, to perform the roll-outs. To ensure reproducibility, all experiments share the same master random seed. For each experiment, we illustrate the progression of the average reward stored in the replay buffer during the training phase. As the replay buffer functions on a first-in, first-out basis, this average provides a dependable measure of the system’s learning advancement from a reward standpoint. The goal of the DRL system is to maximize the reward obtained, irrespective of the selected reward function.
For the QoS and FQoS rewards, we set $\epsilon_{i}^{D}=0.15$ and $\epsilon_{i}^{U}=2$ for all users, with $\mu=2$. For the FQoS reward, we employ $\alpha=\frac{2}{3}$.
To compare the different systems derived from various reward functions, we also show the evolution of rewards obtained with each reward function. The reward function used for the optimization process is termed the decisive reward function, while the others are referred to as informative. By examining the curves for each system, we can effectively compare their performances and behavior using a consistent standard. Additional metrics recorded and stored during the training process, along with the precise code used to generate these metrics.
@@ V-B Downlink and Uplink Power Patterns
Figure 3 shows the downlink and uplink power patterns generated using our simulation framework. This figure presents an example of beam patterns for both downlink and uplink scenarios in a duplex communication system with a single user equipment. The yellow curves represent the beam power patterns transmitted by the BS toward the RIS for the user in the downlink scenario. The purple curves depict the beamforming power patterns at the RIS level using the projected polar profile. It is directed toward either the user or the BS, depending on whether it is downlink or uplink communication. The green rectangle represents an obstacle blocking the initial signal from the BS, highlighting the necessity of using the RIS.
@@ V-C Simulation Results
Figure 4 presents the average reward within the replay buffer at time $t$. This metric serves as an indicator of the system’s learning progress and its improvement concerning decisive reward functions. The first-in-first-out nature of the replay buffer facilitates the analysis of learning trajectories for each system. However, this figure does not allow a direct comparison of reward system performance due to the different scales of the three possible reward functions. It does, however, allow for the monitoring of individual system evolution over time. It is evident that all RIS-controllers utilizing the baseline and QoS rewards demonstrate learning over time, achieving significantly higher rewards compared to the initial training episodes. Conversely, those employing the FQoS reward exhibit limited improvement, potentially attributable to sub-optimal hyperparameter selection and reward shaping.
To enable a comparative assessment of algorithm performance, Figure 5 illustrates the progression of the baseline reward, cf. Equation (4), during the training phase for all systems, despite its relevance to only certain configurations. From this perspective, the systems that demonstrate superior performance, without considering fairness, are those trained with the QoS, cf. Equation (5), and the baseline rewards. Notably, the system trained with the QoS reward and employing DDPG surpasses all other trained systems by at least 15%, achieving a stabilized mean SSR of approximately 8 bps/Hz by the end of training. In comparison, the baseline reward reaches up to 7 bps/Hz for the TD3 variant. Systems trained with FQoS, cf. Equation (10), exhibit inferior performance, achieving only 6 bps/Hz, indicating challenges in balancing performance with fairness considerations.
@@ V-D Discussion
For an evaluation of the RIS-controllers fairness, Figure 6 displays the evolution of the JFI throughout the training process. This index is calculated using Equation (9) for the baseline reward, reflecting the actual signal received by users. The figure reveals that the RIS-controllers previously identified as achieving the highest SSR, specifically those using the QoS reward with DDPG, exhibit lower fairness indices, ranging between 0.64 and 0.61 on average. Given that the corresponding JFI ranges from $\frac{1}{k}$ in the most unfair scenario to 1 in a perfectly fair distribution, these values suggest that one user predominantly receives the majority of the rate. This occurs despite the minimum quality of service constraints being met in the case of the controller using the QoS reward with DDPG.
Systems with lower performance achieve higher average fairness indices, between 0.67 and 0.65. However, this marginal increase in fairness does not compensate for the loss in of pure SSR in Figure 5. To further understand the fairness extremes of the trained RIS-controllers, it is necessary to examine the JFI corresponding to the maximum decisive reward obtained in each episode.
Figure 6: Mean Jain Fairness Index (JFI) for each episode. Smoothed version with a rolling window of size $50$.
Figure 7 shows the JFI obtained for the maximum reward value reached during the episode. It reveals more pronounced behavioral differences among the systems. Those achieving higher performances with respect to the baseline reward exhibit significantly lower fairness, below 0.6, whereas systems utilizing the FQoS maintain a consistent fairness level as previously described. We can conclude that, on average and at peak performance, the current systems are unfair. They heavily favor one user while only satisfying the minimum quality of service constraints for the other. This analysis underscores the importance of embedding fairness signals in reward systems. Although the proposed FQoS reward achieves suboptimal performance and is not without flaws, the study demonstrates that alternative rewards do not offer superior fairness. There is a clear need for extensive hyperparameter research with the existing reward functions, as well as a more comprehensive exploration of strategies to balance signal maximization, quality of service constraints, and fairness. Since we only changed the reward functions, changing the composition of the state could also leverage additional performance and represent an other research direction. This paves the way for enhanced RIS-controllers and ultimately lead to better and fairer B5G communication systems.
Figure 7: Jain Fairness Index (JFI) for the best reward obtained at each episode. Smoothed version with a rolling window of size $50$.
@@ VI Conclusion
We explored the concept of fair communications with respect to DRL-enabled RIS environments. We assumed that RISs must ensure that UE units receive their signals with adequate strength, without other UE being deprived of service due to insufficient power. We investigated such a problem and explored the fairness properties of previous work. We proposed a novel method that aims at obtaining an efficient and fair duplex DRL-RIS system, for multiple legitimate UE units. We reported experimental work and simulation results. We have also released our code and experiments to foster further research on this topic. Future research directions may include an investigation into hyper-parameter optimization and reward engineering to achieve a better balance between pure capacity and perfect fairness, thereby attaining optimal performance. Subsequent studies could concentrate on the detailed analysis of more complex environments by incorporating user mobility patterns and expanding the scale of the system, specifically in terms of the number of users or RIS elements.